Grade 10Mechanics

A particle starsmoving inpositive xdirection with initial velocity of magnitude 10ms-1 with a uniform acceleration of magnitude2ms-2 but directed innegative x direction.What is the distance traversed by the particle in 12 second?

kavita
8 Years agoGrade 10

2 Answers

Khimraj
8 Years ago
time when veocity become zero 
0 = 10 – 2t
t = 5
so distance travelled in first five sec
d1 = 10*5 – ½*2*52 = 25m
distance travelled in last 7 sec
d2 = ½*2*72 = 49m
so total distance travelled = d1 + d2 = 74m
Hope it clears.
yashveer  singh
8 Years ago
yashveer singh 
class 11th
v(0)=10m/s
uniform acc..=-2m/s2
dv / dt=-2
dv=-2dt
∫dv=-2∫dt
v=-2t+c    ----eqn(1)
At t=0,v=10m/s----
10=-2*0+c
c=10
putting the value of c in eqn(1).....
v=-2t+10  --------It is the v-t eqn
now let’s find the time when v=0,
from v-t eqn-----
0=-2t+10
t=5sec  (At t=5sec v=0)
now finding the s-t eqn from the v-t eqn-----
by integration.....
v=-2t+10
ds / dt=-2t+10
ds=-2tdt+10dt
∫ds=-2∫tdt+10∫dt
s=-t2+10t   --------Its the s-t eqn
now finding displacement in t=5sec when v=0,
s=-(5)2+10*5   (from s-t eqn)
s=25m  
let say this displacement as D1
D1=25m
…........................................
…........................................
v(5)=0
uniform acc..=-2m/s2
finding  v-t eqn.........
dv / dt=-2
dv=-2dt
by integation 
v=-2t+c
At t=5,v=0;
v=-2t+10
ds / dt=-2t+10
ds=-2tdt+10dt
now by integrating...
s=-t2+10t
at t=12’displkacement=?,
s=-(12)2+10*12
s=-144+120
s=-24
lets say this displacement as D2
D2=-24m
now total distance=|D1|+|D2 |
25+24=49m
now total displacement=D1+D2
25+(-24)=1m
HOPE THIS ANSWER CLEARS TJHE CONCEPT OF SOLVING THE QUESTION BY DIFFERENTIATION AND INTEGRATION.........................................